OpenStudy (anonymous):

Solve 2cos^2x+cosx-1=0 for the interval [0, 2π)

OpenStudy (mathstudent55):

This is a quadratic equation in \(\cos x\).

OpenStudy (imstuck):

You have to factor this just like you would a "regular" polynomial.

OpenStudy (imstuck):

You get nice and easy factors too! Any ideas?

OpenStudy (anonymous):

I have no idea what to do.

OpenStudy (mathstudent55):

To make it easier to understand, substitute \(\cos x\) with \(u\). Then solve the quadratic equation for \(u\). Then substitute back \(\cos x \) for \(u\), and solve the trig equations.

OpenStudy (mathstudent55):

The above is an idea of what to do. Do you know how to solve a simple quadratic equation by factoring?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

Help?

OpenStudy (mathstudent55):

Do you know how to factor a trinomial, such as 2x^2 + x - 1 ?